# Through the Dual-Prism: A Spectral Perspective on Graph Data Augmentation for Graph Classification

### v1

@@ 4.2 Empirical Evidence
Different importance of high- and low-frequency parts. We then conduct experiments on group-level eigenvalues, i.e., the high- and low-frequency eigenvalues, to gain a broader view of the influence exerted by varying frequency eigenvalues. We introduce noise to eigenvalues across various hyperparameter combinations. Concretely, we use the standard deviation $\sigma$ to determine the magnitude of the noise. The frequency ratio $r_{f}$ dictates the number of eigenvalues $N_{a}$ we might change, while the augmentation probability $r_{a}$ specifies the final eigenvalues sampled for modification. The eigenvalues are arranged in ascending order. The setting of ‘Low’ means that we select candidates from the first $N_{a}$ eigenvalues, while ‘High’ denotes selection from the last $N_{a}$ eigenvalues. As shown in Figure 5b, the orange lines consistently outperform the green lines across all three plots, indicating that the performance associated with perturbing high-frequency eigenvalues consistently exceeds that of their low-frequency counterparts. Moreover, as the frequency ratio $r_{f}$ increases, the accuracy of the ‘Low’ scenario remains relatively stable and low. Contrastingly, for the ‘High’ scenario, a notable decline in accuracy begins once the ratio exceeds around 30%. This suggests that the eigenvalues outside the top 30% of the high-frequency range may start to include more critical information beneficial for graph classification tasks that should not be distorted.
4.3 Theoretical Backing and Insights
Figure 5: (a) Training loss of GIN model on REDDIT-BINARY when graphs are augmented by masking different eigenvalues. (b) Graph classification performance on IMDB-BINARY under various hyperparameters. The lines represent the average accuracy, while the shaded areas indicates the error margins.
@@ 4.3 Theoretical Backing and Insights
The eigenvalues of the Laplacian matrix provide significant insights into various graph properties, as established in prior research (Chung, 1997). Such insights have driven and backed our proposal to modify the high-frequency eigenvalues while preserving their low-frequency counterparts. For example, the second-smallest eigenvalue $\lambda_{1}$, often termed the Fiedler value, quantifies the graph’s algebraic connectivity. A greater Fiedler value indicates a better-connected graph and it is greater than 0 if and only if the graph is a connected graph. The number of times 0 appears as an eigenvalue is the number of connected components in the graph. In addition to the connectivity, the diameter of a graph is also highly related to the eigenvalues – it can be upper and lower bounded from its spectrum (Chung, 1997): $4/n\lambda_{1}\leq d\leq 2[\sqrt{2m/\lambda_{1}}\log_{2}n]$, where $n$ and $m$ denotes the number of nodes and the maximum degree of the graph, respectively. In addition to these widely-used properties, other properties are also highly related to spectrum, including graph diffusion distance (Hammond et al., 2013). In essence, eigenvalues serve as powerful spectral signatures comprising a myriad of structural and functional aspects of graphs.
5 Experiments
Experimental Setup. We evaluate our augmentation method for graph classification tasks under four different settings, including supervised learning, semi-supervised learning, unsupervised learning, and transfer learning. We conduct our experiments on 21 real-world datasets across three different domains, including bio-informatics, molecule, and social network, from the TUDatasets benchmark (Morris et al., 2020), OGB benchmark (Hu et al., 2020a) and ZINC chemical molecule dataset (Hu et al., 2020b). The details of the datasets, baselines, and experimental settings can be found in Appendices A, B, and C, respectively. More empirical results, including the evaluations on hyperparameter sensitivity analysis, can be found in Appendix D.
Table 1: Performance comparisons with GCN and GIN in the supervised learning setting. The best and second best results are highlighted with bold and underline, respectively. * and ** denote the improvement over the second best baseline is statistically significant at level 0.1 and 0.05, respectively(Newey & West, 1987). Baseline results are taken from Ling et al. (2023); Han et al. (2022).
@@ 5 Experiments
[…]
DropEdge 73.20$\pm$5.62 49.00$\pm$2.94 85.15$\pm$2.81 51.19$\pm$1.74 47.08$\pm$0.55 71.61$\pm$4.28 68.32$\pm$1.60 -
DropNode 73.80$\pm$5.71 50.00$\pm$4.85 83.65$\pm$3.63 47.71$\pm$1.75 47.93$\pm$0.64 72.69$\pm$3.55 70.73$\pm$2.02 -
Subgraph 70.90$\pm$5.07 49.80$\pm$3.43 68.41$\pm$2.57 47.31$\pm$5.23 47.49$\pm$0.93 67.93$\pm$3.24 65.05$\pm$4.36 -
[…]
@@
[…]
GIN Vanilla 71.30$\pm$4.36 48.80$\pm$2.54 89.15$\pm$2.47 53.17$\pm$2.26 50.23$\pm$0.83 68.28$\pm$2.47 79.08$\pm$2.12 -
DropEdge 70.50$\pm$3.80 48.73$\pm$4.08 87.45$\pm$3.91 54.11$\pm$1.94 49.77$\pm$0.76 68.01$\pm$3.22 76.47$\pm$2.34 -
DropNode 72.00$\pm$6.97 45.67$\pm$2.59 88.60$\pm$2.52 53.97$\pm$2.11 49.95$\pm$1.70 69.64$\pm$2.98 74.60$\pm$2.12 -
Subgraph 70.40$\pm$4.98 43.74$\pm$5.74 76.80$\pm$3.87 50.09$\pm$4.94 49.67$\pm$0.90 66.67$\pm$3.10 60.17$\pm$2.33 -
@@
[…]
5.1 Supervised learning
Performance. We first evaluate our proposed methods in the supervised learning setting. Following the prior works (Han et al., 2022; Ling et al., 2023), we use GIN and GCN as backbones for graph classification on eight different datasets. Table 1 shows the performance of our proposed methods compared with seven state-of-art (SOTA) baselines, including DropEdge(Rong et al., 2019), DropNode(Feng et al., 2020), Subgraph(You et al., 2020), M-Mixup (Verma et al., 2019), SubMix(Yoo et al., 2022), G-Mixup(Han et al., 2022) and S-Mixup (Ling et al., 2023). According to the results, our DP-Noise method consistently outperforms other existing methods across the majority of datasets, establishing its dominance in effectiveness. DP-Mask also shines, often securing a noteworthy second-place standing. GIN tends to obtain superior outcomes, especially when combined with DP-Noise, which is exemplified by its 61.67% classification accuracy on IMDB-M. Note that on REDD-B, GIN achieves more satisfactory performance than GCN, which is a consistent pattern across baselines but becomes particularly pronounced with our methods. This phenomenon may be attributed to the intrinsic characteristics of GIN and GCN. GIN is known for its precision in delineating complex structural intricacies of graphs (Xu et al., 2018), while GCN is characterized by its smoothing effect (Defferrard et al., 2016). Our methods’ superiority in diversifying the graphs’ structures naturally amplifies GIN’s strengths. In contrast, GCN may not be as adept at leveraging the enhancements offered by our augmentation techniques.
Generalization. Figures 6a and  6b display the test loss and accuracy curves for the IMDB-B dataset, comparing four distinct augmentation strategies: G-mixup, DP-Noise, DP-Mask, and a scenario without any augmentation (i.e., Vanilla). A notable trend is the consistently lower and more stable test loss curves for DP-Noise and DP-Mask in comparison to Vanilla and G-mixup. Concurrently, the accuracy achieved with DP-Noise and DP-Mask is higher. This indicates our proposed methods’ superior generalization and the capacity for enhancing model stability.
@@ 5.1 Supervised learning
5.2 Semi-supervised learning
Performance. We then evaluate our methods in a semi-supervised setting comparing with five baselines, including training from scratch without and with augmentations (denoted as Vanilla and Aug., respectively), GAE (Kipf & Welling, 2016b), Informax (Veličković et al., 2018) and GraphCL(You et al., 2020). Table 2 provides the performance comparison when utilizing 1% and 10% label ratios. At the more challenging 1% label ratio, our DP-Noise achieves SOTA results across all three datasets, and DP-Mask secures the second-best performance on two out of the three datasets. As the label ratio increases to 10%, DP-Noise maintains its efficacy, showcasing excellent performance on six out of seven datasets.
Table 2: Performance comparisons in the semi-supervised learning setting. The best and second best results are highlighted with bold and underline, respectively. 1% or 10% is the label ratio. The metric is accuracy (%). * and ** denote the improvement over the second best baseline is statistically significant at level 0.1 and 0.05, respectively. Baseline results are taken from You et al. (2020).
@@ 5.2 Semi-supervised learning
1% Vallina 60.72$\pm$0.45 - - 57.46$\pm$0.25 - - 54.25$\pm$0.22
1% Aug. 60.49$\pm$0.46 - - 58.40$\pm$0.97 - - 56.36$\pm$0.42
1% GAE 61.63$\pm$0.84 - - 63.20$\pm$0.67 - - 59.44$\pm$0.44
[…]
1% GraphCL 62.55$\pm$0.86 - - 64.57$\pm$1.15 - - 58.56$\pm$0.59
1% DP-Noise 63.43$\pm$1.39 - - 65.94$\pm$3.13 - - 60.06$\pm$2.72
[…]
10% Vallina 73.72$\pm$0.24 70.40$\pm$1.54 73.56$\pm$0.41 73.71$\pm$0.27 86.63$\pm$0.27 51.33$\pm$0.44 60.87$\pm$0.17
10% Aug. 73.59$\pm$0.32 70.29$\pm$0.64 74.30$\pm$0.81 74.19$\pm$0.13 87.74$\pm$0.39 52.01$\pm$0.20 60.91$\pm$0.32
[…]
10% Infomax 74.86$\pm$0.26 72.27$\pm$0.40 75.78$\pm$0.34 73.76$\pm$0.29 88.66$\pm$0.95 53.61$\pm$0.31 65.21$\pm$0.88
[…]
Augmentation Pairing Efficiency. To investigate optimal combinations that could potentially enhance performance, we evaluate the synergistic effects on accuracy gain (%) when pairing DP-Noise and DP-Mask with different augmentation methods using the same setting in You et al. (2020). From Figures 6c and  6d, overall, the diverse accuracy gains across datasets indicate that there is no one-size-fits-all “partner” for DP-Noise and DP-Mask; the efficacy of each combination varies depending on the dataset. However, generally, both DP-Noise and DP-Mask exhibit enhanced performance when paired with dropN at a large degree.
5.3 Unsupervised representation learning
We next evaluate our strategies in the unsupervised learning setting and compare them with 12 baselines, including three graph kernel methods (GL (Pržulj, 2007), WL(Shervashidze et al., 2011) and DGK(Yanardag & Vishwanathan, 2015)), four representation learning methods (node2vec (Grover & Leskovec, 2016), sub2vec(Adhikari et al., 2018), graph2vec(Narayanan et al., 2017) and InfoGraph(Sun et al., 2019)) and five GCL-based methods (GraphCL(You et al., 2020), MVGRL(Hassani & Khasahmadi, 2020), AD-GCL(Suresh et al., 2021), JOAO(You et al., 2021) and GCL-SPAN(Lin et al., 2022)). Table 3 shows the performance of our methods in an unsupervised setting. From the results, DP-Noise and DP-Mask surpass other baselines on five out of seven datasets. Notably, compared with another spectral-based method GCL-SPAN (Lin et al., 2022), our methods outperform it on most datasets, especially on the molecules dataset NCI1 (increase of around 11.5% accuracy). This can be explained by that despite GCL-SPAN is also spectral-based, it in fact modifies the spatial graph while optimizing in the spectral realm. In contrast, our methods directly make alterations in the spectral domain to preserve the structural information. Given the critical role that structural patterns in molecular data play on classification tasks, the enhanced performance underscores the efficiency of our direct spectral modifications in creating more effective and insightful augmented graphs for graph classification.
@@ 5.3 Unsupervised representation learning
Table 3: Performance comparisons in the unsupervised learning results. The best and second best results are highlighted with bold and underline, respectively. The metric is accuracy (%). * and ** denote the improvement over the second best baseline is statistically significant at level 0.1 and 0.05, respectively.
[…]
5.4 Transfer learning
We lastly conduct transfer learning experiments on molecular property prediction in the manner of Hu et al. (2020b) to evaluate the capability of our methods for learning generalizable encoders. Specifically, we initially pre-train models on the extensive chemical molecule dataset ZINC (Sterling & Irwin, 2015), then fine-tune the models on eight distinct datasets within a similar domain. We draw comparisons between our methods and six baselines, including a reference model without pre-training (referred to No-Pre-Train), Informax (Veličković et al., 2018), EdgePred (Hamilton et al., 2017), AttrMasking (Hu et al., 2020b), ContextPred (Hu et al., 2020b) and GraphCL You et al. (2020). The comparative results are shown in Table 4. Our methods demonstrated SOTA performance, outperforming competitors on half of the datasets. Especially, our methods consistently outperform the conventional GraphCL method, which indicates our data augmentation methods as better choices for graph contrastive learning. Notably, DP-Mask achieves an 83.52% ROC-AUC score on ClinTox, exceeding the performance of GraphCL by a substantial margin (nearly 10%). These findings demonstrate the enhanced efficacy of our techniques in the transfer learning setting for graph classification tasks.
@@ 5.4 Transfer learning
Table 4: Performance comparisons in the transfer learning setting. The best and second best results are highlighted with bold and underline, respectively. The metric is ROC-AUC scores (%). * and ** denote the improvement over the second best baseline is statistically significant at level 0.1 and 0.05, respectively.
[…]
6 Conclusion

### v3

@@ A Spectral Lens on Graph Data Augmentations
Different importance of high- and low-frequency parts. We then conduct experiments on group-level eigenvalues, i.e., the high- and low-frequency ones, to gain a broader view of the influence exerted by varying frequency eigenvalues. We introduce noise to eigenvalues across various hyperparameter combinations. Concretely, we use the standard deviation $\sigma$ to determine the magnitude of the noise. The frequency ratio $r_{f}$ dictates the number of eigenvalues $N_{a}$ we might change, while the augmentation probability $r_{a}$ specifies the final eigenvalues sampled for modification. The eigenvalues are arranged in ascending order. The setting of ’Low’ means that we select candidates from the first $N_{a}$ eigenvalues, while ’High’ denotes selection from the last $N_{a}$ eigenvalues. As shown in Figure 3b, the orange lines consistently outperform the green lines across all three plots, indicating that the performance associated with perturbing high-frequency eigenvalues consistently exceeds that of their low-frequency counterparts. Moreover, as the frequency ratio $r_{f}$ increases, the accuracy of the ’Low’ scenario remains relatively stable and low. Contrastingly, for the ’High’ scenario, a notable decline in accuracy begins once the ratio exceeds around 30%. This suggests that the eigenvalues outside the top 30% of the high-frequency range may start to include more critical information beneficial for graph classification tasks that should not be distorted.
Theoretical Backing and Insights
@@
The eigenvalues of the Laplacian matrix provide significant insights into various graph properties (Chung 1997). Such insights have driven and backed our proposal to modify the high-frequency eigenvalues while preserving their low-frequency counterparts. For example, the second-smallest eigenvalue $\lambda_{1}$, often termed the Fiedler value, quantifies the graph’s algebraic connectivity. A greater Fiedler value indicates a better-connected graph and it is greater than 0 if and only if the graph is connected. The number of times 0 appears as an eigenvalue is the number of connected components in the graph. In addition to the connectivity, the diameter of a graph is also highly related to the eigenvalues – it can be upper and lower bounded from its spectrum (Chung 1997): $4/n\lambda_{1}\leq d\leq 2[\sqrt{2m/\lambda_{1}}\log_{2}n]$, where $n$ and $m$ denotes the number of nodes and the maximum degree of the graph. Besides these widely-used ones, other properties are also highly related to spectrum, e.g., graph diffusion distance (Hammond, Gur, and Johnson 2013). In essence, eigenvalues serve as powerful spectral signatures comprising a myriad of structural and functional aspects of graphs.
@@
GCN Vanilla 72.80$\pm$4.08 49.47$\pm$2.60 84.85$\pm$2.42 49.99$\pm$1.37 46.90$\pm$0.73 71.43$\pm$2.60 72.38$\pm$1.45 96.80$\pm$0.06
DropEdge 73.20$\pm$5.62 49.00$\pm$2.94 85.15$\pm$2.81 51.19$\pm$1.74 47.08$\pm$0.55 71.61$\pm$4.28 68.32$\pm$1.60 96.53$\pm$0.07
DropNode 73.80$\pm$5.71 50.00$\pm$4.85 83.65$\pm$3.63 47.71$\pm$1.75 47.93$\pm$0.64 72.69$\pm$3.55 70.73$\pm$2.02 96.52$\pm$0.14
Subgraph 70.90$\pm$5.07 49.80$\pm$3.43 68.41$\pm$2.57 47.31$\pm$5.23 47.49$\pm$0.93 67.93$\pm$3.24 65.05$\pm$4.36 96.49$\pm$0.17
M-Mixup 72.00$\pm$5.66 49.73$\pm$2.67 87.05$\pm$2.47 51.49$\pm$2.00 46.92$\pm$1.05 71.16$\pm$2.87 71.58$\pm$1.79 -
@@
DP-Mask 76.00$\pm$3.62 51.20$\pm$1.73 76.70$\pm$1.34 52.42$\pm$2.78 47.25$\pm$1.12 73.60$\pm$3.10 62.45$\pm$3.80 96.90$\pm$0.24*
GIN Vanilla 71.30$\pm$4.36 48.80$\pm$2.54 89.15$\pm$2.47 53.17$\pm$2.26 50.23$\pm$0.83 68.28$\pm$2.47 79.08$\pm$2.12 96.54$\pm$0.11
DropEdge 70.50$\pm$3.80 48.73$\pm$4.08 87.45$\pm$3.91 54.11$\pm$1.94 49.77$\pm$0.76 68.01$\pm$3.22 76.47$\pm$2.34 96.48$\pm$0.09
DropNode 72.00$\pm$6.97 45.67$\pm$2.59 88.60$\pm$2.52 53.97$\pm$2.11 49.95$\pm$1.70 69.64$\pm$2.98 74.60$\pm$2.12 96.51$\pm$0.16
Subgraph 70.40$\pm$4.98 43.74$\pm$5.74 76.80$\pm$3.87 50.09$\pm$4.94 49.67$\pm$0.90 66.67$\pm$3.10 60.17$\pm$2.33 96.49$\pm$0.13
@@
DP-Mask 76.30$\pm$2.56 51.60$\pm$1.32 93.25$\pm$1.19** 56.50$\pm$0.80* 49.11$\pm$1.30 72.79$\pm$1.95** 80.30$\pm$2.01 96.98$\pm$0.29
Table 1: Performance comparisons with GCN and GIN in the supervised learning setting. The best and second best results are highlighted with bold and underline, respectively. * and ** denote the improvement over the second best baseline is statistically significant at level 0.1 and 0.05 (Newey and West 1987). Baseline results are taken from (Ling et al. 2023; Han et al. 2022).
Dataset NCI1 COLLAB GITHUB
Vallina 60.72$\pm$0.45 57.46$\pm$0.25 54.25$\pm$0.22
Aug. 60.49$\pm$0.46 58.40$\pm$0.97 56.36$\pm$0.42
GAE 61.63$\pm$0.84 63.20$\pm$0.67 59.44$\pm$0.44
Infomax 62.72$\pm$0.65 61.70$\pm$0.77 58.99$\pm$0.50
GraphCL 62.55$\pm$0.86 64.57$\pm$1.15 58.56$\pm$0.59
DP-Noise 63.43$\pm$1.39 65.94$\pm$3.13 60.06$\pm$2.72
DP-Mask 62.43$\pm$1.08 65.68$\pm$1.66* 59.70$\pm$0.53
Table 2: Performance comparisons in the semi-supervised learning setting with 1% label ratio. The best and second best results are highlighted with bold and underline, respectively. Baseline results are taken from (You et al. 2020).
Experiments
Experimental Setup. We evaluate our method for graph classification under four different settings, including supervised learning, semi-supervised learning, unsupervised learning, and transfer learning. We conduct our experiments on 21 real-world datasets across three different domains, including bio-informatics, molecule, and social network, from the TUDatasets (Morris et al. 2020), OGB (Hu et al. 2020a) and ZINC chemical molecule dataset (Hu et al. 2020b).
@@
Vallina 73.72$\pm$0.24 70.40$\pm$1.54 73.56$\pm$0.41 73.71$\pm$0.27 86.63$\pm$0.27 51.33$\pm$0.44 60.87$\pm$0.17
Aug. 73.59$\pm$0.32 70.29$\pm$0.64 74.30$\pm$0.81 74.19$\pm$0.13 87.74$\pm$0.39 52.01$\pm$0.20 60.91$\pm$0.32
GAE 74.36$\pm$0.24 70.51$\pm$0.17 74.54$\pm$0.68 75.09$\pm$0.19 87.69$\pm$0.40 53.58$\pm$0.13 63.89$\pm$0.52
Infomax 74.86$\pm$0.26 72.27$\pm$0.40 75.78$\pm$0.34 73.76$\pm$0.29 88.66$\pm$0.95 53.61$\pm$0.31 65.21$\pm$0.88
GraphCL 74.63$\pm$0.25 74.17$\pm$0.34 76.17$\pm$1.37 74.23$\pm$0.21 89.11$\pm$0.19 52.55$\pm$0.45 65.81$\pm$0.79
DP-Noise 75.30$\pm$0.58** 74.73$\pm$1.01 76.91$\pm$0.81 77.05$\pm$0.82** 89.38$\pm$0.95 54.45$\pm$0.64** 65.59$\pm$0.88
DP-Mask 74.88$\pm$1.84 71.37$\pm$4.18 75.64$\pm$0.81 76.90$\pm$0.62** 88.62$\pm$0.63 52.80$\pm$0.59 64.95$\pm$1.03
@@
Table 3: Performance comparisons in the semi-supervised learning setting with 10% label ratio. The best and second best results are highlighted with bold and underline, respectively. The metric is accuracy (%). * and ** denote the improvement over the second-best baseline is statistically significant at levels 0.1 and 0.05. Baseline results are taken from (You et al. 2020).
Dataset NCI1 PROTEINS DD MUTAG REDD-B REDD-M5 IMDB-B
sub2vec 52.84$\pm$1.47 $5$ 61.05$\pm$15.80 71.48$\pm$0.41 36.68$\pm$0.42 55.26$\pm$1.54
graph2vec 73.22$\pm$1.81 73.30$\pm$2.05 - 83.15$\pm$9.25 75.78$\pm$1.03 47.86$\pm$0.26 71.10$\pm$0.54
InfoGraph 76.20$\pm$1.06 74.44$\pm$0.31 75.23$\pm$0.39 89.01$\pm$1.13 82.50$\pm$1.42 53.46$\pm$1.03 73.03$\pm$0.87
GraphCL 77.87$\pm$0.41 74.39$\pm$0.45 78.62$\pm$0.40 86.80$\pm$1.34 89.53$\pm$0.84 55.99$\pm$0.28 71.14$\pm$0.44
MVGRL 68.68$\pm$0.42 74.02$\pm$0.32 75.20$\pm$0.55 89.24$\pm$1.31 81.20$\pm$0.69 51.87$\pm$0.65 71.84$\pm$0.78
AD-GCL 69.67$\pm$0.51 73.59$\pm$0.65 74.49$\pm$0.52 89.25$\pm$1.45 85.52$\pm$0.79 53.00$\pm$0.82 71.57$\pm$1.01
JOAO 72.99$\pm$0.75 71.25$\pm$0.85 66.91$\pm$1.75 85.20$\pm$1.64 78.35$\pm$1.38 45.57$\pm$2.86 71.60$\pm$0.86
GCL-SPAN 71.43$\pm$0.49 75.78$\pm$0.41 75.78$\pm$0.52 89.12$\pm$0.76 83.62$\pm$0.64 54.10$\pm$0.49 73.65$\pm$0.69
DP-Noise 79.69$\pm$0.70 74.60$\pm$0.43 78.59$\pm$0.23 87.63$\pm$1.98 90.90$\pm$0.32** 55.54$\pm$0.15 71.42$\pm$0.41
DP-Mask 79.47$\pm$0.22 74.70$\pm$0.29 79.97$\pm$1.09** 89.98$\pm$1.36 91.21$\pm$0.24** 55.92$\pm$0.49 71.78$\pm$0.37
@@
Table 4: Performance comparisons in the unsupervised learning results. The best and second best results are highlighted with bold and underline. The metric is accuracy (%). * and ** denote the improvement over the second-best baseline is statistically significant at levels 0.1 and 0.05, respectively. Baseline results are taken from (You et al. 2020; Lin, Chen, and Wang 2022).
Supervised Learning
Performance. We first evaluate our proposed methods in the supervised learning setting. Following the prior works (Han et al. 2022; Ling et al. 2023), we use GIN and GCN as backbones for graph classification on eight different datasets. Table 1 shows the performance of DP methods compared with seven state-of-art (SOTA) baselines, including DropEdge(Rong et al. 2019), DropNode(Feng et al. 2020), Subgraph(You et al. 2020), M-Mixup (Verma et al. 2019), SubMix(Yoo, Shim, and Kang 2022), G-Mixup(Han et al. 2022) and S-Mixup (Ling et al. 2023). According to the results, DP-Noise consistently outperforms other existing methods across the majority of datasets, establishing its dominance in effectiveness. DP-Mask also shines, often securing a second-place standing. GIN tends to obtain superior outcomes, especially when combined with DP-Noise, exemplified by its 61.67% accuracy on IMDB-M. Note that on REDD-B, GIN achieves more satisfactory performance than GCN, which is a consistent pattern across baselines but becomes particularly pronounced with our methods. This phenomenon may be attributed to the intrinsic characteristics of GIN and GCN: GIN is known for its precision in capturing the complex structures of graphs (Xu et al. 2018), while GCN is characterized by its smoothing effect (Defferrard, Bresson, and Vandergheynst 2016). Our methods’ superiority in diversifying the graphs’ structures naturally amplifies GIN’s strengths. In contrast, GCN may not be as adept at leveraging the enhancements offered by our techniques.
Generalization. Figures 6a and  6b display the test loss and accuracy curves for the IMDB-B dataset, comparing four distinct augmentation strategies: G-mixup, DP-Noise, DP-Mask, and a scenario without any augmentation (i.e., Vanilla). The consistently lower and more stable test loss curves for DP-Noise and DP-Mask is comparison to Vanilla and G-mixup. Concurrently, the accuracy achieved with DP-Noise and DP-Mask is higher, indicating the DP methods’ superior generalization and the capacity for enhancing model stability.
Semi-supervised Learning
Performance. We then evaluate our methods in a semi-supervised setting comparing with five baselines, including training from scratch without and with augmentations (denoted as Vanilla and Aug.), GAE (Kipf and Welling 2016b), Informax (Veličković et al. 2018) and GraphCL(You et al. 2020). Tables 2 and  3 provide the performance comparison when utilizing 1% and 10% label ratios. At the more challenging 1% label ratio, DP-Noise achieves SOTA results across all three datasets, and DP-Mask secures the second-best performance on two out of the three datasets. As the label ratio increases to 10%, DP-Noise maintains its efficacy, showing SOTA performance on six out of seven datasets.
Augmentation Pairing Efficiency. To investigate optimal combinations that could potentially enhance performance, we evaluate the synergistic effects on accuracy gain (%) when pairing DP-Noise and DP-Mask with different augmentation methods using the same setting in (You et al. 2020). From Figures 6c and  6d, overall, the diverse accuracy gains across datasets indicate that there is no one-size-fits-all “partner” for DP-Noise and DP-Mask; the efficacy of each combination varies depending on the dataset. However, generally, both DP-Noise and DP-Mask exhibit enhanced performance when paired with dropN at a large degree.
Unsupervised Representation Learning
We next evaluate DP methods in unsupervised learning and compare them with eight baselines, including three representation learning methods (sub2vec(Adhikari et al. 2018), graph2vec(Narayanan et al. 2017) and InfoGraph(Sun et al. 2019)) and five GCL-based methods (GraphCL(You et al. 2020), MVGRL(Hassani and Khasahmadi 2020), AD-GCL(Suresh et al. 2021), JOAO(You et al. 2021) and GCL-SPAN(Lin, Chen, and Wang 2022)). Table 4 shows the performance of our methods in an unsupervised setting. From the results, DP methods surpass other baselines on five out of seven datasets. Notably, compared with another spectral-based method GCL-SPAN (Lin, Chen, and Wang 2022), our methods outperform it on most datasets, especially on the molecules dataset NCI1 (an increase of around 11.5% accuracy). This superiority can be attributed to the fact that while GCL-SPAN is also spectral-based, it operates on graph modification in the spatial domain rather than directly manipulating the spectral domain like DP methods. Given the critical role that structural patterns in molecular data play on classification tasks, the enhanced performance underscores the efficiency of our direct spectral modifications in creating more effective and insightful augmented graphs.
Transfer Learning
We conduct transfer learning experiments on molecular property prediction in the manner of (Hu et al. 2020b). Specifically, we initially pre-train models on the extensive chemical molecule dataset ZINC (Sterling and Irwin 2015), then fine-tune the models on eight distinct datasets within a similar domain. Due to space constraints, the table containing performance comparisons and analysis is provided in the Appendix.
Conclusion